# STONE: Pioneering the One-to-N Backdoor Threat in 3D Point Cloud

### v2

@@ 3.4.2 Identification of Key Parameters
After selecting the spherical trigger as the base entity, the next critical step is to identify which parameters to vary for realizing the one-to-N attack. The chosen parameters must satisfy two key requirements: first, any variation must preserve the trigger’s stealthiness to avoid detection; second, the parameters must significantly influence the backdoored model’s behavior to ensure discriminative power across different target classes.
In our approach, we deliberately exclude radius variation as a discriminative parameter, as size changes would significantly compromise imperceptibility and violate the first requirement. Instead, we focus exclusively on spatial position—specifically, the sphere’s center coordinates—as our primary parameterization method. This allows us to maintain a fixed minimal radius across all trigger instances while leveraging spatial diversity to create distinct configurations.
To rigorously validate that spatial position indeed governs backdoor efficacy—thus satisfying the second requirement—we employ two complementary approaches to establish this causal relationship: theoretical analysis through the Neural Tangent Kernel framework and empirical investigation through systematic experimentation.
Theoretical Analysis. Our theoretical foundation builds upon the Neural Tangent Kernel (NTK) framework. Lemma 1 formally establishes that for an insertion-based backdoor attack, the probability of successful backdoor activation satisfies $f_{\bm{\theta}}(\bm{X}^{\prime}_{R_{1}})[t_{i}]<f_{\bm{\theta}}(\bm{X}^{\prime}_{R_{0}})[t_{i}]$ when an identical spherical trigger is implanted at two spatially separated regions $R_{0}$ and $R_{1}$, where $\bm{X}^{\prime}_{R_{0}}$ and $\bm{X}^{\prime}_{R_{1}}$ denote poisoned samples with triggers at positions $R_{0}$ and $R_{1}$ respectively, both targeting class $t_{i}$, and $R_{0}$ represents the trigger position used during the training phase. This result demonstrates that backdoor activation depends not only on the presence of the trigger but also critically on its spatial position.
Building upon this foundation, Lemma 2 further reveals through NTK analysis that the predictive probability for the target class exhibits a distance-dependent relationship: $f_{\bm{\theta}}(\bm{X}^{\prime}_{R})[t_{i}]$ decreases as the Euclidean distance between $R$ and $R_{0}$ increases. This theoretical insight directly informs our trigger placement strategy. When two triggers are placed too close together, the probability of misclassification between their respective target classes becomes significant, leading to mutual interference that compromises the discriminative power of our one-to-N framework. Therefore, maintaining adequate spatial separation between triggers is essential to ensure reliable target class assignment.
@@ 3.4.3 Trigger Configuration
We implement two concrete configurations within the STONE framework:
@@ 3.4.4 Parameter Set Optimization
The core objective of parameter set optimization is to select optimal configurations for $N$ triggers within the normalized 3D bounding box, with spatial location serving as the discriminative parameter. To this end, maintaining sufficient separation between triggers is critical for preserving the spatial specificity, as established in our theoretical analysis. This requirement leads to an NP-hard maximin optimization problem aimed at maximizing the minimum pairwise Euclidean distance. We address this challenge by developing efficient greedy algorithms that iteratively select trigger positions from a candidate set to ensure adequate separation. (See Appendix Algorithms 1 and 2 for detailed pseudocode.)
@@ 3.4.5 End-to-End Attack Pipeline
Model Training. The model is trained following Eq. (4) using poisoned samples from all $N$ targets in a single process, learning the association between spherical configurations and target labels. The resulting model exhibits dual behavior: correct classification of clean inputs and targeted misclassification when triggers are present. This establishes the complete one-to-N capability, enabling selective activation as shown in Fig. 2, where triggers at $(0.95,0.95,0.95)$ and $(0.05,0.05,0.05)$ are mapped to classes Girl and Table, respectively.
Multi-Target Activation. During inference, multi-target activation is accomplished by implanting the pre-configured spherical triggers into clean inputs. The activation mechanism is illustrated in Fig. 1(b) and (c) where, for instance, a single sphere at $(0.95,0.95,0.95)$ leads to a misclassification to girl in a model poisoned with single-sphere triggers, and a dual-sphere pair at $(0.95,0.95,0.95)$ and $(0.95,0.95,0.05)$ results in a misclassification to girl in a model poisoned with dual-sphere triggers.
@@ 4.2.1 Empirical Evidence of Spatial Specificity
This overall decreasing trend provides the foundational principle for trigger placement in the STONE framework. The existence of such a relationship implies that, for any chosen ASR threshold defining acceptable interference, there exists a corresponding critical distance. In our specific experimental configuration, an ASR threshold of 5% corresponds to an observed critical distance of approximately 0.26. This empirical observation demonstrates that to prevent significant interference between any two triggers in the one-to-N framework, their separation must exceed the critical distance specific to the model and data distribution. Therefore, our objective of maximizing the minimum inter-trigger distance is directly motivated by the need to mitigate interference between triggers, thereby ensuring reliable discrimination among all target classes in the STONE framework..
@@ 4.2.2 Impact of Poisoning Ratio
Performance vs. Poisoning Ratio: The experimental results demonstrate a clear positive correlation between poisoning ratio and attack success rate across all configurations. As illustrated in Fig. 4, the average ASR exhibits a characteristic saturation pattern, with rapid improvement in the low poisoning ratio regime (0.3%-2%) followed by by converging to perfect success (ASR = 1.0) at higher ratios (3%-4%).
@@
ModelNet10 exhibits intermediate performance, achieving perfect ASR (100%) at 3% poisoning for single-sphere and maintaining this level for dual-sphere triggers. ModelNet40, with its greater class diversity (40 categories), shows slightly lower susceptibility, requiring approximately 4% poisoning to achieve performance close to 100%. This performance hierarchy, with ShapeNetPart outperforming ModelNet10, which in turn outperforms ModelNet40, suggests that dataset complexity and class diversity influence attack difficulty, though the performance differences remain practically marginal for operational deployment.
@@ 5 Conclusion
This work establishes STONE, the first comprehensive framework that demonstrates how to realize one-to-N backdoor attacks in 3D point clouds using configurable spherical triggers. The framework’s feasibility is theoretically grounded in Neural Tangent Kernel analysis, which shows that poisoned models can learn to associate distinct spatial trigger configurations with different target labels. This theoretical foundation is extensively validated through rigorous experiments across multiple datasets and architectures, confirming that multi-target backdoor attacks represent a severe and adaptable threat to 3D vision systems.
Our work provides both a theoretical basis and a practical implementation for this new attack paradigm, establishing spherical triggers as an effective mechanism for dynamic target control in 3D domains. The STONE framework opens a new research direction by proving the viability of spatial configuration as a discriminative mechanism, paving the way for developing more flexible, stealthy, and defense-resistant one-to-N attacks.
@@ Appendix A Theoretical Analysis of Spatial Specificity and Spatial Sensitivity
A.1 Theoretical Foundations
Lemma 1 (Spatial Specificity of Spherical Triggers).
Lemma 2 (Spatial Sensitivity).
A.2 Key Assumptions and Simplifications
@@
A.3 Detailed Proofs
A.3.1 Proof of Lemma 1

### v3

@@ 3.4.2 Identification of Key Parameters
After selecting the sphere as the base trigger design, the critical step is to identify which parameters to vary for the one-to-N attack. The chosen parameters must satisfy two requirements: first, any variation must preserve the trigger’s stealthiness; second, the parameters must significantly influence the backdoored model’s behavior to ensure discriminative power across target classes with low inter-target interference.
We deliberately exclude radius variation as a discriminative parameter, as size changes would compromise imperceptibility. Instead, we focus exclusively on spatial position—the sphere’s center coordinates—as our primary parameterization. This maintains a fixed minimal radius for stealth while leveraging spatial diversity to create distinct trigger configurations.
To rigorously validate that spatial position governs backdoor efficacy—thus satisfying the second requirement—we establish this causal relationship through two complementary approaches: theoretical analysis via the Neural Tangent Kernel framework and empirical investigation through systematic experimentation.
Theoretical Analysis. Our theoretical foundation builds upon the Neural Tangent Kernel (NTK) framework, leading to the following lemmas:
Lemma 1 (Spatial Specificity of Spherical Triggers).
Lemma 2 (Spatial Sensitivity).
Lemma 1 formally establishes that for an insertion-based backdoor attack, the probability of successful backdoor activation satisfies $f_{\bm{\theta}}(\bm{X}^{\prime}_{R_{1}})[t_{i}]<f_{\bm{\theta}}(\bm{X}^{\prime}_{R_{0}})[t_{i}]$ when an identical spherical trigger is implanted at two spatially separated regions $R_{0}$ and $R_{1}$, where $\bm{X}^{\prime}_{R_{0}}$ and $\bm{X}^{\prime}_{R_{1}}$ denote poisoned samples with triggers at positions $R_{0}$ and $R_{1}$ respectively, both targeting class $t_{i}$, and $R_{0}$ represents the trigger position used during the training phase. This result demonstrates that backdoor activation depends not only on the presence of the trigger but also critically on its spatial position.
Lemma 2 further reveals through NTK analysis that the predictive probability for the target class exhibits a distance-dependent relationship: $f_{\bm{\theta}}(\bm{X}^{\prime}_{R})[t_{i}]$ decreases exponentially as the Euclidean distance between $R$ and $R_{0}$ increases. This theoretical insight directly informs our trigger placement strategy. When two triggers are placed too closely, the interference between their target class mappings becomes significant, compromising discriminative power. Therefore, it crucial to maintain sufficient spatial separation ensures reliable target assignment. On the other hand, because the interference decays exponentially with distance, multiple triggers can be deployed within a normalized point cloud, enabling a practical one-to-N attack framework.
@@ 3.4.3 Trigger Configuration
We implement two concrete configurations within the STONE method:
@@ 3.4.4 Parameter Set Optimization
The core objective of parameter set optimization is to select optimal configurations for $N$ triggers within the normalized 3D bounding box, with spatial location serving as the discriminative parameter. To this end, maintaining sufficient separation between triggers is critical for minimizing inter-target interference, as established in our theoretical analysis. This requirement leads to an NP-hard maximin optimization problem aimed at maximizing the minimum pairwise Euclidean distance. We address this challenge by developing greedy algorithms that iteratively select trigger positions from a candidate set to ensure adequate separation. (See Appendix Algorithms 1 and 2 for detailed pseudocode.)
Figure 3: End-to-end pipeline of the STONE framework (single-sphere design). The framework (1) implants class-specific spherical triggers into clean samples; (2) trains a model to associate trigger configurations with target classes while maintaining accuracy on clean inputs; (3) enables multi-target activation during inference, where applying a specific trigger causes misclassification to its designated target, as illustrated in Fig.2(b)(c).
@@ 3.4.5 End-to-End Attack Pipeline
Model Training. The model is trained following Eq. (4) using poisoned samples from all $N$ targets in a single process, learning the association between spherical configurations and target labels. The resulting model exhibits dual behavior: correct classification of clean inputs and targeted misclassification when triggers are present.
Multi-Target Activation. During inference, multi-target activation is accomplished by implanting the pre-configured spherical triggers into clean inputs.
@@ 4.2.1 Empirical Evidence of Spatial Specificity
This overall decreasing trend provides the foundational principle for trigger placement in the STONE framework. The existence of such a relationship implies that, for any chosen ASR threshold defining acceptable interference, there exists a corresponding critical distance. In our specific experimental configuration, an ASR interference threshold of 5% corresponds to an observed critical distance of approximately 0.26. This empirical observation demonstrates that to prevent significant interference between any two triggers in the one-to-N framework, their separation must exceed the critical distance specific to the model and data distribution. Therefore, our objective of maximizing the minimum inter-trigger distance is directly motivated by the need to mitigate interference between triggers, thereby ensuring reliable discrimination among all target classes in the STONE framework.
@@ 4.2.2 Impact of Poisoning Ratio
Performance vs. Poisoning Ratio: The experimental results demonstrate a clear positive correlation between poisoning ratio and attack success rate across all configurations. As illustrated in Fig. 5, the average ASR exhibits a characteristic saturation pattern, with rapid improvement in the low poisoning ratio regime (0.3%-2%) followed by by converging to perfect success at higher ratios (3%-4%).
@@
ModelNet10 exhibits intermediate performance, achieving perfect at 3% poisoning for single-sphere and maintaining this level for dual-sphere triggers. ModelNet40, with its greater class diversity (40 categories), shows slightly lower susceptibility, requiring approximately 4% poisoning to achieve performance close to 100%. This performance hierarchy, with ShapeNetPart outperforming ModelNet10, which in turn outperforms ModelNet40, suggests that dataset complexity and class diversity influence attack difficulty, though the performance differences remain practically marginal for operational deployment.
@@ 5 Conclusion
This work establishes STONE, the first approach that realizes the one-to-N universal backdoor attack in 3D point clouds under the black box and dirty label setting—a critical threat model previously unexplored. The feasibility of this attack is supported by a key theoretical insight: the interference between distinct target mappings exhibits an exponential decay with respect to the distance between their corresponding trigger configurations in the parameter space. This property allows for the design of multiple, minimally interfering triggers within a confined spatial domain. This theoretical foundation is extensively validated through rigorous experiments across multiple datasets and network architectures, confirming that multi-target backdoor attacks represent a severe and practically viable threat to 3D vision systems.
Our work thus bridges the critical research gap in one-to-N backdoor universal attacks for 3D point clouds, transforming this concept from an intuitive possibility into a practically realized and theoretically substantiated threat. By proving the viability of parametric configuration as a robust discriminative mechanism, the STONE framework opens a new research direction and paves the way for developing more flexible, stealthy, and defense-resistant backdoor strategies in 3D domains.
@@ Appendix A Theoretical Analysis of Spatial Specificity and Spatial Sensitivity
A.1 Key Assumptions and Simplifications
@@
A.2 Detailed Proofs
A.2.1 Proof of Lemma 1